OpenStudy (anonymous):

What is period T of x(t) x(t) = cos(3t) + cos(10t) I was thinking that T = 2*pi, is this correct?

OpenStudy (jamesj):

The period of the function is the smallest number T > 0 such that x(t) = x(t+T) Remember that cos a + cos b = 2 cos ( (a+b)/2 ) . cos( (a-b)/2 )

OpenStudy (anonymous):

13pi/10

OpenStudy (anonymous):

hmm if i type it in wolframalpha i also got 2pi :) http://www.wolframalpha.com/input/?i=period+cos%283t%29+%2B+cos%2810t%29

OpenStudy (anonymous):

but jamesj, what about Cn? i can write x(t) as: \[1/2e ^{j3t}+1/2e ^{-j3t}+1/2e ^{j10t}+1/2e ^{-j10t}\] Is \[C1=1/2 + 1/2\] can i just add those 1 terms together?

OpenStudy (jamesj):

2pi is not the answer. 2pi is the period of cos x and sin x. Let me ask you, what the period of cos(3x) all be itself?

OpenStudy (anonymous):

thats \[2\pi/3 = T\]

OpenStudy (anonymous):

period of (cos3t )=2pi/3,n (cos10t)=2pi/10 use the trick take LCM of periods

OpenStudy (jamesj):

right. Now, as x(t) = cos(3t) + cos(10t) = 2 . cos (13t/2) . cos(7t/2) what period works for both expressions?

OpenStudy (anonymous):

but i dont know that answer of: 2 . cos (13t/2) . cos(7t/2)

OpenStudy (jamesj):

cos(13t/2) has period 2pi/(13/2) = 4pi/13 cos(7t/2) has period 2pi/(7/2) = 4pi/7 Now find their lcm. Or, as Sam suggests, find the lcm of period of (cos3t ) = 2pi/3 and period(cos10t) = 2pi/10

OpenStudy (jamesj):

If you're having trouble with the lcm concept for periods, work some examples: what is period of - cos x + cos 2x - cos 2x + cos 3x

OpenStudy (anonymous):

but what is lcm? and how can i calculate it :(

OpenStudy (jamesj):

LCM = lowest common multiple. Take integer multiples of both periods until you find the the smallest pair which are equal. For example, for cos x + cos 2x the periods are 2pi and pi. Clearly the LCM of those two periods is 2pi. Hence the period of cos x + cos 2x is 2pi. For cos 2x + cos 3x the periods are pi and 2pi/3 Now the integer multiples of pi are: pi, 2pi, 3pi, ... intger multiples of 2pi/3 are: 2pi/3, 4pi/3, 2pi, ... hence their LCM is 2pi and the period of cos 2x + cos 3x is 2pi. Use the same method here.

OpenStudy (anonymous):

oke, so T1 = 2pi/3-> 2pi/3, 4pi/3, 6pi/3.... T2 = 2pi/10 -> 2pi/10, 4pi/10, 6pi/10....

OpenStudy (anonymous):

until i find the match?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

there must be a faster way

OpenStudy (anonymous):

divide T1 with T2?

OpenStudy (jamesj):

not necessarily. you're in university now; don't expect the problems to be so elementary.

OpenStudy (anonymous):

hmm.. i was thinking about dividing the 2 periods

OpenStudy (jamesj):

That doesn't work, most emphatically not.

OpenStudy (jamesj):

Just do it, it's not hard.

OpenStudy (anonymous):

ahhh alright.. and my other question... about Cn? have you read that one?

OpenStudy (jamesj):

What is your other question exactly?

OpenStudy (anonymous):

i have to find Cn of x(t) (fourier) You can write x(t) as form of Euler, as i typed in above... Cn for n=1, is 1/2 + 1/2, but can i add there 2 terms together...

OpenStudy (jamesj):

I don't understand your question. You're trying to find the Fourier series of something? Your function x(t) = cos(3t) + cos(10t) is a Fourier series in its own right. [ x(t) = 0 cos t + 0 cos(2t) + 1 cos(3t) + 0 cos(4t) + ... etc. ] But what is it exactly you're asking?

OpenStudy (anonymous):

a formula to find cn is: \[cn*e ^{jn \omega t}\] like cos(t) = (e^j3t + e^-j3t)/2 for n=+/-1 cn = 1/2 for other values its zero

OpenStudy (anonymous):

the same is for cos(10t) for n=+/- 1 Cn = 1/2 other n Cn = 0

OpenStudy (anonymous):

But what is C1 for x(t)

OpenStudy (jamesj):

Wait. Are you trying to find the coefficients c_n such that \[ x(t) = \sum_{i=-\infty}^\infty c_n e^{njx} \] ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

i just gave you the answers, but can i add my 2 answers?

OpenStudy (jamesj):

Then notice that if x(t) = cos(3x) and nothing else, then c_n = 0 except for n = +3, -3. Likewise for cos(10x), except for n = +10, -10 Hence for x(t) = cos(3x) + cos(10x), the only non-zero coefficients c_n are those for which c_n = -10, -3, 3 and 10

OpenStudy (anonymous):

ahh oke... but what is C_n for -10, -3, 3 and 10?

OpenStudy (jamesj):

c'mon! What is it if x(t) = cos(3x) and nothing else? What is \[ c_3, c_{-3} \]

OpenStudy (anonymous):

thats 1/2

OpenStudy (jamesj):

yes.

OpenStudy (anonymous):

but what is x(t) = cos(3t) + cos(10t) OWHHH pellet! >.< its 1/2 for n=-10,-3, 3 and 10

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

i was just confused.. about adding 1/2 with the other 1/2

OpenStudy (anonymous):

buttt.. all clear now :) thank you